OpenStudy (anonymous):

if x+a=(b/3)x and b cannot be equal to 3, the x=

OpenStudy (anonymous):

drawing it will be helpful, thanks

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x+a= \frac{b}{3}x \LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

And \(\normalsize\color{blue}{ b≠3 \LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x+a= \frac{b}{3}x \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ x+a= \frac{bx}{3} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ x- \frac{bx}{3}+a=0 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \frac{3x}{3}- \frac{bx}{3}+a=0 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \frac{3x-bx}{3}+a=0 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \frac{(3-b)x}{3}+a=0 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \frac{(3-b)x}{3}=-a \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \frac{x}{3}=\frac{-a}{(3-b)} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ x=\frac{-3a}{(3-b)} \LARGE\color{white}{ \rm │ }}\) this is why the say that b is not 3, because the ddnominator is going to be a zero if it will be 3.

OpenStudy (anonymous):

oh, now i remenber , thank you

OpenStudy (solomonzelman):

Anytime !

OpenStudy (anonymous):

oh i have another equation question

OpenStudy (anonymous):

4√3 5√3=

OpenStudy (anonymous):

i keep getting 20√3, but the answer should be 20√3^9

OpenStudy (anonymous):

i would like to know how?

OpenStudy (solomonzelman):

By saying \(\normalsize\color{blue}{ 4\sqrt{3}~5\sqrt{3} \LARGE\color{white}{ \rm │ }}\) they likely mean, \(\normalsize\color{blue}{ 4\times \sqrt{3}\times~5\times\sqrt{3} \LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

So, to calculate this, \(\normalsize\color{blue}{ 4\times \sqrt{3}\times~5\times\sqrt{3} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 4\times~5\times\sqrt{3}\times \sqrt{3} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 20\times\sqrt{3}^{~2}\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 20\times3\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 60\LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

the equation is in this picture

OpenStudy (anonymous):